### v1

# Simultaneous Gesture Classification and Localization with an Automatic Gesture Annotation Model

Abstract
Training a real-time gesture recognition model heavily relies on annotated data. However, manual data annotation is costly and demands substantial human effort. In order to address this challenge, we propose a novel annotation model that can automatically annotate gesture classes and identify their temporal ranges. Our ablation study demonstrates that our annotation model design surpasses the baseline in terms of both gesture classification accuracy (3-4% improvement) and localization accuracy (71-75% improvement). We believe that this annotation model has immense potential to improve the training of downstream gesture recognition models using unlabeled datasets.
Index Terms—  Gestural input, machine learning, annotation model
1 Introduction
Hand gesture recognition technology is prevalent in fields like augmented reality (AR) and virtual reality (VR). Different types of hand gesture data, such as RGB, depth, and 3D skeletons, are used for recognition. With the rise of AR and VR head-mounted devices (HMDs), real-time 3D hand skeleton data is accessible and consistent across various devices and sensors [1]. In this paper, we focus on gesture classification and localization using skeleton data.
Building a gesture recognition model generally involves data collection, annotation, and training. Annotation encompasses classifying gestures and determining their temporal range. As deep-learning techniques dominate the state-of-the-art gesture recognition models [2], there’s a substantial need for labeled data. Insufficient data could lead to overfitting in deep neural networks [3]. The challenge is that data annotation, especially detailed frame-wise annotation, is time-intensive and expensive.
To combat data sparsity in gesture recognition, researchers have used data augmentation techniques like Generative Adversarial Networks for gesture data transformation [3] and unrealistic distortions like cutout [4] and mixup [5]. Approaches like few-shot learning have also been explored to reduce dataset size needs [6]. However, the reliance on existing annotated datasets and the associated annotation costs remain significant concerns. The potential of leveraging large unlabeled datasets and the automation of the annotation process remains largely untapped.
We introduce an annotation model that can predict unlabeled gesture data classification and localization. This model allows for large-scale, high-quality annotations without significant human intervention, serving as a foundation for real-time gesture recognition systems.
Using the SHREC’2021 [7] and Online DHG [8] datasets, we conducted thorough tests on our model. An ablation study revealed our model’s superior performance over the baseline model. To our knowledge, this is the first attempt at presenting an automatic model for simultaneous gesture classification and localization.
2 Related Work
Traditionally, manual annotation software has been employed to annotate gestures, with human annotators manually detecting the start and end points of each gesture. However, this process is time-consuming and requires significant labor. As a result, there has been growing interest in developing automatic annotation tools. While various fields have explored this idea (e.g., [9, 10, 11]), relatively fewer works have focused on gesture data specifically. One relevant work by Kratz et al. [12] presents preliminary research on automatically segmenting motion gestures tracked by IMUs. They suggest that by recognizing gesture execution phases from motion data, it might be possible to automatically identify the start and end points of gestures. However, their work primarily focuses on gesture localization and does not address gesture classification. More closely related to our work, Lenaga et al.[13] propose the use of an Active Learning (AL)[14] framework to automatically detect sign language gesture occurrences in RGB videos. Their approach requires manual annotation for only a small subset of the videos, benefiting researchers studying multimodal communication. However, their primary application area is sign language gestures in RGB videos, which predominantly focuses on spatial data. In contrast, our model focuses on temporal data for gesture annotation.
3 Problem Formulation
This section formulates gesture annotation and gesture recognition and discusses their difference. We also explain why simply using a conventional gesture recognition model does not suit data annotation and why an annotation model is necessary.
It is important to clarify that when referring to gesture recognition, we specifically focus on real-time gesture recognition. In this context, real-time gesture recognition involves identifying gestures within a realistic sequence that includes multiple gesture classes as well as background activities (no gestures), commonly observed in a sensor stream.
To achieve real-time gesture recognition, a typical approach is to utilize a sliding-window-based technique. At each sliding step, a window is fed into the model, producing class-conditional probabilities. The predicted gesture class within the window is determined by taking the maximum value from the probabilities. However, the conventional real-time gesture recognition model is not well-suited for data annotation, particularly for gesture nucleus localization. Post-processing techniques need to be employed to localize the gesture nucleus. However, these post-processing steps often rely on heuristic methods (e.g., [1]), which may be susceptible to variations in gestures and contextual differences.
Furthermore, although both gesture annotation and real-time gesture recognition aim for accurate gesture classification, they have different objectives. While real-time gesture recognition models prioritize detecting gestures as quickly as possible, gesture annotation additionally focuses on identifying the nucleus of the gesture as accurately as possible. In contrast, an annotation model can incorporate gesture nucleus localization as an inherent training objective since it does not need to optimize for recognition latency. This enables better architectural design choices to improve the performance of the annotation task.
4 Proposed Model
In this section, we describe our model for annotating gesture data.
4.1 Design Considerations
Fig. 1: Due to the extended input window of the annotation model, the true label has the capacity to encompass multiple gestures simultaneously. In the illustrated case, the true labels are {1, - , 3, - ,2}, where “-” represents background activities (no gesture). The output of the annotation model is a sequence of predicted labels that has the same length as inputs. The loss function for the proposed annotation model is a Connectionist Temporal Classification (CTC) loss. The backbone of the model is an A-ResLSTM we adopted from Shen et al. [1].
We propose an annotation model optimized for gesture classification and localization and address four key design considerations in our architecture.
  1. 1.
    Small vs Large Window Size: Conventionally, small windows ensure accurate classification by focusing on a single gesture. Although small windows lead to low recognition latency, they have limited temporal context. Hence, we utilize a larger window size with the sliding-window technique to process multiple gestures concurrently.
  2. 2.
    One Prediction vs Multiple Predictions Per Frame: Standard classification models provide a single vector of class-conditional probabilities per frame. This many-to-one mapping can lose fine-grained temporal information. Conversely, our many-to-many architecture retains richer temporal information, enabling dynamic adjustments between windows, reducing prediction noise.
  3. 3.
    Strongly-Segmented vs Weakly-Segmented Training Data: While models trained using cross-entropy loss need strongly-segmented data, obtaining such data is labor-intensive. We employ Connectionist Temporal Classification (CTC) loss, aligning unsegmented inputs with outputs without requiring strong segmentation, a technique found effective in seq2seq models.
  4. 4.
    Heuristic Parameters vs Automation: Traditional models like Shen et al. [1] employ heuristic threshold parameters to pinpoint gesture locations. As these parameters can affect localization, we favor a model devoid of heuristic parameters. Leveraging CTC loss, we directly localize the gesture nucleus by identifying the spike position in the CTC’s output without post-processing.
4.2 Annotation Model
After considering these four design factos, we now introduce our model’s individual components in detail.
4.2.1 Seq2seq Model
We design a seq2seq model that is a many2many model. It has an input window $\mathbf{W}$ of frame length $L$ (window size) and an output which is a matrix of probabilities $\mathbf{M}\in\mathbb{R}^{L\times(K+1)}$. It defines the probabilities of detecting a gesture (or no gesture) $k$ at time $t$ in an input window $\mathbf{W}$: $p(k,t\mid\mathbf{W})=\mathbf{M}^{t}_{k}\forall t\in\left[0,N\right)$. The backbone of our seq2seq model is an A-ResLSTM model we adopted from Shen et al. [1], which is one of the state-of-the-art neural network architectures for gesture recognition. This network is composed of residual blocks [15] and bi-directional LSTM layers [16] with an attention layer [17].
4.2.2 Connectionist temporal classification (CTC)
Fig. 2: Demonstration of the network output probabilities from a CTC and CE (cross-entropy) trained network versus the ground truth [18]. The spike of CTC loss clearly captures the gesture nucleus, while the curve of CE loss needs post-processing. For example, the CTC loss successfully distinguishes two consecutive same gestures, while the CE loss confuses them together.
Compared to conventional cross-entropy (CE) loss, CTC loss has the benefit of producing output with distinct, sharp predictions that can accurately capture consecutive gestures. In contrast, a model trained with CE tends to blend predictions of the same class together (see Figure 2). This means that in order to correctly identify when one gesture has been performed, the output from a CE-trained classification model must be processed further with some heuristic methods with errors (e.g., a single-time activation algorithm [1]). However, with CTC, the sharp nature of the predictions automatically eliminates such error-prone step [18, 19, 20].
4.2.3 Dynamic Adjustment
Due to the sliding-window fashion, for a window $\mathbf{W}$ of length $L$ sliding at step $N$, the model outputs $\mathbf{M}$ for each window. $\mathbf{M}$ consists of $L$ vectors of class-conditional probabilities p. Therefore, the model will produce $L/N$ times class-conditional probabilities p at each step/frame, forming a sequence $(\textbf{p}_{1},\textbf{p}_{2},...,\textbf{p}_{L/N})$. We perform a dynamic adjustment on labeling and nucleus localization. Specifically, we average these predictions to form one vector of class-conditional probabilities for each step, i.e. $\textbf{p}_{average}=\text{average}(\textbf{p}_{1},\textbf{p}_{2},...,\textbf{p}_{L/N})$ We determine the location of the gesture’s nucleus by identifying the position of the highest point in the gesture’s prediction.
4.2.4 Decoding
We use greedy search for the dynamic adjustment output, which approximates the solution by taking the most likely class at each step in the matrix $\mathbf{M}$. It is efficient to compute as it simply concatenates the most active outputs at each step.
5 Experiments and Analysis
SHREC’2021 [7] Online DHG [8]
Static one - four, OK, menu NA
Dynamic left, right, circle, v, cross rotation (counter-)clockwise, right/left, up/down, cross, plus, v, shake
Fine Dynamic grab, pinch, tab, deny, knob, expand grab, pinch, tab, expand
Size 180 seqs of 3-5 gestures occurring sequentially 280 seqs of 10 gestures occurring sequentially
Table 1: Contents of the datasets. Static gestures are characterized by keeping a fixed hand pose for a minimum amount of time, while dynamic gestures are characterized by a single trajectory with an unchanged hand pose or with finger articulation over time. Fine dynamic gestures are characterized by a single trajectory with changing hand pose.
SHREC’2021 [7] Online DHG [8]
Accuracy NNLE Accuracy NNLE
A-ResLSTM [1] w/ CE Loss (Baseline Model) 88.3 ($\pm$ 2.34) 0.42 ($\pm$ 0.07) 89.8 ($\pm$ 3.12) 0.44 ($\pm$ 0.03)
A-ResLSTM w/ CTC loss 90.6 ($\pm$ 1.45) 0.17 ($\pm$ 0.03) 91.5 ($\pm$ 3.35) 0.16 ($\pm$ 0.04)
A-ResLSTM w/ CTC loss & many2many 91.9 ($\pm$ 1.12) 0.14 ($\pm$ 0.03) 92.7 ($\pm$ 3.35) 0.15 ($\pm$ 0.04)
A-ResLSTM w/ CTC loss & many2many & dynamic adjustment 92.6 ($\pm$ 1.58) 0.12 ($\pm$ 0.02) 93.2 ($\pm$ 1.76) 0.11 ($\pm$ 0.03)
Table 2: Ablation study results by adding our multiple design parts step by step. NNLE is our metric for gesture localization (Normalized Nucleus Labeling Error). The numbers in the brackets indicate std.
We conducted a series of experiments to evaluate our proposed annotation model.
5.1 Evaluation Datasets
In order to ensure a thorough evaluation of our Model, we perform our experiments on two publicly available datasets, SHREC’2021 [7] and Online DHG [8]. These datasets consist of sequences of unsegmented gesture data, encompassing various types of gestures. The specific details of the two datasets are as follows: Both datasets already provided a predefined split between training and testing data. Thus, we conducted evaluations on the designated testing sets.
5.2 Evaluation Metrics
We introduce our metrics to evaluate the tasks of gesture classification and localization.
  1. 1.
    Gesture Classification Accuracy: Gesture annotation would generate a sequence of classification output. We use the Levenshtein distance (also known as minimum edit distance) to evaluate the gesture classification performance. It is a metric defined as the minimum number of single-character (in our case, classification output of each frame) insertions, deletions, and substitutions required to transform one string into another. The accuracy for recognition performance is defined as:
    $Accuracy=1-\frac{\operatorname{levenshtein}(y_{predict},y_{true})}{\operatorname{length}(y_{true})}$ (1)
    where $y_{predict}$ and $y_{true}$ are the predicted and true list of labels of the gestures, respectively.
  2. 2.
    Nucleus Localization Error: We propose a metric Normalized Nucleus Localization Error (NNLE) to measure how accurately our model can locate the gesture nucleus. When a gesture is recognized, we define the time for the start/end of the gesture as $idx_{start}$/$idx_{end}$, and the detected location of gesture nucleus as $idx_{nucleus}$. For a accurate nucleus localization, $idx_{start}\leq idx_{nucleus}\leq idx_{end}$. NNLE is defined as:
    $NNLE=\frac{idx_{activation}-(idx_{start}+idx_{end})/2+1}{idx_{end}-idx_{start}+1}$ (2)
    A smaller NNLE means our annotation model is more accurate for gesture localization (i.e., the nucleus is closer to the center of the gesture).
5.3 Training Details
Our implementation was based on TensorFlow 2. We utilized the Adam optimizer with a starting learning rate of 0.0001. To prevent overfitting, we employed early stopping with a patience of 5. For both the Online DHG and SHREC’2021 datasets, we used a window length of 200 for our model. It is important to note that the length of an individual gesture within the datasets ranged from 20 to 50 frames, corresponding to a duration of approximately 0.4 to 1 second.
5.4 Ablation Study
Our novel annotation model essentially consists of multiple design blocks: (1) using the CTC loss instead of a basic CE loss, (2) using a many-to-many architecture (i.e., seq2seq) instead of a many-to-one architecture, and (3) using dynamic adjustment on labeling and nucleus localization. To evaluate the effectiveness of each design block, we performed an ablation study through a step-by-step addition process.
Table 2 illustrates the ablation study results. Compared to the baseline (A-ResLSTM trained with Cross-Entropy (CE) loss), our model significantly reduces the gesture localization error by 0.3 (71.4%) and 0.33 (75.0%). Regarding gesture classification accuracy, our model design achieves improvements of 4.3% on the SHREC’2021 dataset and 3.4% on the Online DHG dataset compared to the baseline.
Each of the three blocks contributes approximately equally to the overall gesture classification accuracy improvement. For gesture nucleus localization, the main contribution stems from the application of the Connectionist Temporal Classification (CTC) loss. This highlights the unique character of the CTC-trained model in producing outputs with a spiky nature, eliminating the need for complex post-processing techniques (see Figure 2). In contrast, CE-trained networks require post-processing to determine the final label, often relying on heuristic methods.
6 Conclusion
This paper presents a model that enables accurate and automated annotation of gesture data, encompassing gesture classification and localization. The ablation study reveals that this novel annotation model design surpasses the baseline model, achieving a 4.3% higher class labeling accuracy and a 71.4% improvement in nucleus localization accuracy on the SHREC’2021 dataset (3.4% and 75.0% respectively on the Online DHG dataset).